One of the questions on a survey asked adults a series of questions on attitudes toward automation. The results showed that 2,979 out of 4,135 respondents are worried about a future in which robots and computers can do many human jobs; 2,774 are worried about the development of algorithms that can evaluate and hire job candidates; and 2,234 are worried about the development of driverless vehicles. (Round your answers to four decimal places.)

(a) Develop a point estimate of the proportion of respondents who are worried about a future in which robots and computer can do many human jobs.

(b) Develop a point estimate of the proportion of respondents who are worried that the development of algorithms that can evaluate and hire job candidates.

(c) Develop a point estimate of the proportion of respondents who are worried about the development of driverless vehicles.
Answer #1

a)

point estimate of the proportion of respondents who are worried about a future in which robots and computer can do many human jobs. = 2979/4135 = 0.7204


b)

point estimate of the proportion of respondents who are worried that the development of algorithms that can evaluate and hire job candidates = 2774/4135 = 0.6709


c)

point estimate of the proportion of respondents who are worried about the development of driverless vehicles. = 2234/4135 = 0.5403